# Rethinking Data Input for Point Cloud Upsampling

### v1

@@
Abstract
In recent years, point cloud upsampling has been widely applied in fields such as 3D reconstruction and surface generation. However, existing point cloud upsampling inputs are all patch based, and there is no research discussing the differences and principles between point cloud model full input and patch based input. In order to compare with patch based point cloud input, this article proposes a new data input method, which divides the full point cloud model to ensure shape integrity while training PU-GCN. This article was validated on the PU1K and ABC datasets, but the results showed that Patch based performance is better than model based full input i.e. Average Segment input. Therefore, this article explores the data input factors and model modules that affect the upsampling results of point clouds.
Keywords:
@@ Abstract
Deep learning Point cloud Upsampling Data input 3D reconstruction
1 Introduction
@@
The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model. Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision [5, 21, 17, 28].
Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields. In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention [1]. Reverse engineering involves transforming physical objects into digital models for analysis, visualization, and manufacturing. In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software [9], if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless [22]. Therefore, the generation of model point clouds is crucial.
@@
Figure 1: The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
@@
It is beyond dispute that upsampling is a type of generative task. The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net [26] on the basis of pointnet++ [18]. Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud. Moving on to MPU/3PU [25], KNN is used for interpolation and EdgeConv [24] is used to extract features for hierarchical reconstruction through recursive upsampling. Implementing upsampling PU-GAN [12] based on GAN [6] generation architecture, guiding the generator to upsample and approximate the real point cloud. PU-GCN [19] modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results. And Transformer’s point cloud upsampling network [20, 23]. There are also methods to create coarse point clouds and then refine them by combining global and local methods [13].
@@
The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution. The inherent limitations have also been addressed by the work of upsampling point clouds at any scale [27, 8] and capturing local information [10].
@@
Nevertheless, these works still have limitations, which are on the training dataset. Existing point cloud upsampling work [19, 20, 27, 8, 10, 15] without exception used the PU1K dataset [19], which was used based on the methods of MPU and PU-GAN [25, 12]. In order to obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory. Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering. It is necessary to explore a new point cloud input method to avoid the above issues. Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD. In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset [11]. In addition, in order to achieve personalized point cloud model generation in one step after reverse engineering, it is necessary to understand which part of the encoder and decoder, namely feature extraction and data expansion modules, determines the performance of upsampling in point cloud upsampling.
@@
Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method. Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS). After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19] network, and cross validation was performed on both the PU1k and ABC datasets. And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
@@
In order to solve the above problems, this paper proposes a new data input method. Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one. After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19], and cross validation was performed on both the PU1k [19] and ABC datasets [11]. And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
@@
2 Methodology
2.1 Overview
As shown in Fig. 1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale. The reason for choosing PU-GCN as the backbone of the network is that MPU and PU-GCN have similar feature extraction methods. Both use KNN to construct a graph of the input point cloud, with nodes being points and edges defining $k$ nearest point neighborhoods. Then, EdgeConv uses these distances to calculate the edge features of each point. The subsequent PU-Transformer in the upsampling module adopted the Nodeshuffle proposed by PU-GCN, which forms a dense feature map compared to MPU’s Duplicate.
@@ 2.1 Overview
Ergo, in order to avoid memory consumption and computational complexity, the previously proposed patch based end-to-end training has emerged. In situations where resources are limited, dividing the mesh into multiple patches for upsampling is more practical and efficient. But there are also disadvantages to this.
@@
Next, this article will theoretically discuss the differences between the two input methods.
(a) Patch based data input
(b) AS method data input
Figure 2: Schematic diagram of two data input methods for point cloud upsampling
2.2 Data Input Based on Patch
In this case, whichs is showing in Fig. 2(a), it is necessary to divide the Mesh into multiple patches. Assuming the entire Mesh contains $M$ vertices, divide it into $K$ patches. $\mathcal{P}_{k}$ contains $M_{k}$ vertices, where $\sum_{k=1}^{K}M_{k}\geq M$. Convert each patch to a point cloud and perform upsampling. After converting the vertices of each patch to a point cloud, it is assumed that the point cloud contains $N_{k}$ points and we want to upsample them into $N_{k}^{\prime}$ points, typically $N_{k}^{\prime}>N_{k}$.
@@ 2.2 Data Input Based on Patch
2.3 Data Input Based on Average Segment
In this process, like the example in Fig. 2, we can assume that the distribution of sampling points follows a uniform distribution. Assuming that the range of a point cloud is a bounded three-dimensional spatial area, we can use uniform distribution to simulate the sampling process of points.
Specifically, assuming the spatial range of a point cloud is $[x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\times[z_{\text{min}},z_{\text{max}}]$, we can consider the coordinates of each point $(x,y,z)$ as three independent random variables, whose distributions are $\text{Uniform}(x_{\text{min}},x_{\text{max}})$, $\text{Uniform}(y_{\text{min}},y_{\text{max}})$ and $\text{Uniform}(z_{\text{min}},z_{\text{max}})$.
@@ 2.3 Data Input Based on Average Segment
The pseudocode for the above method is showing in Algorithm 1.
@@ 3.1 Settings
  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN [19]. Overall, PU1K consists of 1147 3D models, divided into 1020 training samples and 127 testing samples. The training set includes 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore [4]. The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore. Covering 50 object categories.
The original PU1K dataset was used to adapt to patch based upsampling pipelines, and the training data was generated from patches in the 3D mesh through Poisson sampling. Specifically, the training data includes 69000 samples, each with low resolution: 256 input points and high resolution: 1024 groundtruth points which is four times larger than input. In addition, the method based on AS mentioned in Chapter 2 above also modified the PU1K dataset by uniformly dividing the original 1020 training samples into 8 sub-models that maintain the shape integrity of the model from 2048 points as input in a single model and 8192 points in ground truth. The training data consists of 8160 samples. Similarly, each sample has low resolution: 256 input points and high resolution: 1024 groundtruth points.
@@
  • ABC Dataset: The ABC Dataset [11] is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications. It is currently one of the largest publicly available datasets for geometric deep learning research.
The vast CAD dataset brings diversity, with models covering a wide range of geometric shapes and categories, including architecture, furniture, mechanical parts, and more. And each model is composed of explicitly parameterized curves and surfaces, which can provide real data in areas such as differentiation, crack segmentation, geometric feature detection, and shape reconstruction. This article is based on a subset of ABC 10k in the ABC dataset, which is a CAD model with a quantity of 10000. On this basis, two 4x upsampling experiments were conducted, with the inputs being the complete model and the patch model, respectively. Each sample has low resolution: 512 input points and high resolution: 2048 groundtruth points.
@@
Training details: Straightforwardly, the PU-GCN used in this article, where the network structure is shown in the architecture of (a) in Fig. 3, is implemented using PyTorch. Due to TensorFlow and CUDA versions, this article is unable to re-create the original author’s Python environment in PU-GCN within the TensorFlow deep learning framework. And the experiments were conducted on a single NVIDIA GeForce RTX 3060 LHR GPU running on a Ubuntu 22.04 operating system. In terms of training hyperparameters, this article follows the settings of the original PU-GCN in terms of network structure. Notwithstanding, a batch size of $32$ with $50$ training epochs was used, with an initial learning rate of $5\times 10^{-4}$, a decay rate of $0.05$, and so on. In addition, the loss function also uses Chamfer Distance loss to train PU-GCN, minimizing the average nearest point distance between the input set $\mathcal{P}\in\mathbb{R}^{N\times 3}$ and the output set $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, thus achieving efficient and effective convergence.
@@
Figure 3: The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure. (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
3.2 Evaluation Metrics
@@
During the testing process, this article adopted the common practices used in previous work on point cloud upsampling [19, 20, 27, 8, 10, 15]. That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points. Then, using the PU-GCN model to upsample the subset Patch and AS at a scale of $r$.
@@
In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed. Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points. This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) [7], (ii) Hausdorff distance (HD) [2], and (iii) model parameter quantity. The lower the values under these indicators, the better the upsampling performance.
3.3 Results
@@
PU1K dataset: Tab. 1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset. By upsampling at resolutions of 256-1024 and 2048-8192, it can be seen that the Patch based method outperforms the AS method proposed in this paper in all two metrics of distance measurement.
@@
Table 1: The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
@@
ABC10K dataset: we also conducted point cloud upsampling experiments using the ABC10K dataset. Perform upsampling at a resolution of 512-2048. As shown in Tab. 2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
Table 2: The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
@@
Although the idea, which is originated from overall uniform random sampling, article proposed name Average Segment (AS) method has many advantages in theory, actual experimental results show that the Patch based upsampling method performs better on Chamfer distance (CD) and Hausdorff distance (HD). One possible reason for this may be that when upsampling on a patch, each patch contains fewer point clouds, and the graph convolutional network can extract local features more finely. Due to each patch of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD. Other possibilities that could cause this situation will be discussed in Chapter 4.
3.4 Ablations
Test sets cross validation: This article will conduct cross validation on the PU-GCN trained on the PU1K and ABC10K datasets under the validation sets of both datasets. To understand the impact of the model used for training data on upsampling results. The results are shown in Tab. 3. When the Patch method and AS method are trained on the PU1K dataset and tested on the ABC10K dataset, the CD and HD metrics are significantly superior to those trained and tested on the ABC10K dataset. However, when trained on the ABC10K dataset and tested on the PU1K dataset, there was no significant difference between the CD and HD indicators. It should be noted that in the aforementioned cross validation, the Patch method is still better than the AS method.
Table 3: The results of swapping test sets based on Patch method and AS method on two datasets.
@@ 3.4 Ablations
Architecture Impact: This article breaks down the important modules of PU-GCN, aiming to explore which part has the greatest impact on point cloud upsampling. To this end, the original network backbone has been increased or decreased in three ways: removing DenseGCN (b), adding Refiner (c), and removing DenseGCN while adding Refiner (d), all of them are showed in Fig. 3. The experiment was trained on the PU1K dataset and tested on the test sets of the PU1K dataset and ABC10K dataset. The results are shown in Tab. 4. The results showed that upsampling on the test set with three resolutions of 256-1024, 512-2048 and 2048-8192 in PU1K performed better than the original PU-GCN with DenseGCN, regardless of the presence or absence of Refiner, under the condition of removing DenseGCN. Under the upsampling scale of 2048-8192, the uniform segmentation method that removes DenseGCN is better than the Patch method.
Table 4: Perform ablation experiments on different architectures of PU-GCN using two data input methods.
@@
without Inception DenseGCN Patch PU1K $256\to 1024$ 35.061 54.340 99.152
@@
with Refiner but without Inception DenseGCN Patch PU1K $256\to 1024$ 34.309 54.070 130.945
@@
3.5 Visualizations
Visualize the above experimental results now, which directly displays qualitative comparisons. The qualitative results of upsampling using different point cloud datasets are shown in Fig. 4 and Fig. 5. It can be seen that the input based on Patch performs better in distinguishing edges on the boundaries compared to the AS method. This is particularly evident on the CAD based ABC dataset, as shown in Fig. 5. However, in Fig. 4, it can be seen that the AS method performs better locally in the model. Relatively speaking, it is denser and the connections between points are tighter. In addition, even though DenseGCN performs better on CD and HD on the PU1K dataset, this results in the point cloud lacking local features in visualization. On the other hand, Refiner can solve the problem of identifying local features in the model, achieving the utilization of both global and local features.
Figure 4: Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
Figure 5: Visualization example on the ABC dataset.
4 Discussion and Future Work
4.1 Discussion

### v2

@@
Abstract.
Point cloud upsampling is crucial for tasks like 3D reconstruction. While existing methods rely on patch-based inputs, and there is no research discussing the differences and principles between point cloud model full input and patch based input. Ergo, we propose a novel approach using whole model inputs i.e. Average Segment input. Our experiments on PU1K and ABC datasets reveal that patch-based inputs consistently outperform whole model inputs. To understand this, we will delve into factors in feature extraction, and network architecture that influence upsampling results.
@@ Abstract.
Deep learning, Point cloud, Upsampling, Data input, 3D reconstruction
ccs: Computing methodologies Computer vision
1. Introduction
@@
The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model. Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision (Endres et al., 2013; Roriz et al., 2021; Naseer et al., 2018; Zhou and Tuzel, 2018).
@@
Figure 1. The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields. In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention (Alai, 2013). In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software (Kao and Lin, 1998), if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless (Shatnawi et al., 2017). Therefore, the generation of model point clouds is crucial.
@@
The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net (Yu et al., 2018) on the basis of pointnet++ (Qi et al., 2017). Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud. Moving on to MPU/3PU (Yifan et al., 2019), KNN is used for interpolation and EdgeConv (Wang et al., 2019) is used to extract features for hierarchical reconstruction through recursive upsampling. Implementing upsampling PU-GAN (Li et al., 2019) based on GAN (Goodfellow et al., 2020) generation architecture, guiding the generator to upsample and approximate the real point cloud. PU-GCN (Qian et al., 2021) modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results. And Transformer’s point cloud upsampling network (Qiu et al., 2022; Vaswani et al., 2017). There are also methods to create coarse point clouds and then refine them by combining global and local methods (Li et al., 2021).
@@
The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution. The inherent limitations have also been addressed by the work of upsampling point clouds at any scale (Zhao et al., 2022; He et al., 2023) and capturing local information (Kim et al., 2024).
@@
Nevertheless, these works still have limitations, which are on the training dataset. Existing point cloud upsampling work (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022) without exception used the PU1K dataset (Qian et al., 2021), which was used based on the methods of MPU and PU-GAN (Yifan et al., 2019; Li et al., 2019). To obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory. Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering. It is necessary to explore a new point cloud input method to avoid the above issues. Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD. In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset (Koch et al., 2019).
@@
Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method. Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS). After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021) network, and cross validation was performed on both the PU1k and ABC datasets. And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
@@
To solve the above problems, this paper proposes a new data input method. Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one. After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021), and cross validation was performed on both the PU1k (Qian et al., 2021) and ABC datasets (Koch et al., 2019). And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
@@
2. Methodology
2.1. Overview
As shown in Figure  1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale. The reason for choosing PU-GCN as the backbone of the network is that MPU and PU-GCN have similar feature extraction methods. Both use KNN to construct a graph of the input point cloud, with nodes being points and edges defining $k$ nearest point neighborhoods. Then, EdgeConv uses these distances to calculate the edge features of each point. The subsequent PU-Transformer in the upsampling module adopted the Nodeshuffle proposed by PU-GCN, which forms a dense feature map compared to MPU’s Duplicate.
@@ 2.1. Overview
Ergo, to avoid memory consumption and computational complexity, the previously proposed patch based end-to-end training has emerged. Dividing the mesh into multiple patches for upsampling is more practical and efficient. But there are also disadvantages to this.
@@
2.2. Data Input Based on Patch
In this case, it is necessary to divide the Mesh into multiple patches. Assuming the entire Mesh contains $M$ vertices, divide it into $K$ patches. $\mathcal{P}_{k}$ contains $M_{k}$ vertices, where $\sum_{k=1}^{K}M_{k}\geq M$. Convert each patch to a point cloud and perform upsampling. After converting the vertices of each patch to a point cloud, it is assumed that the point cloud contains $N_{k}$ points and we want to upsample them into $N_{k}^{\prime}$ points, typically $N_{k}^{\prime}>N_{k}$.
@@ 2.2. Data Input Based on Patch
2.3. Data Input Based on Average Segment
We can assume that the distribution of sampling points follows a uniform distribution. Assuming that the range of a point cloud is a bounded three-dimensional spatial area, we can use uniform distribution to simulate the sampling process of points.
@@ 2.3. Data Input Based on Average Segment
The pseudocode for the above method is showing in Algorithm  1.
@@ 3.1. Settings
  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN (Qian et al., 2021). Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore (Chang et al., 2015). The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore. Covering 50 object categories.
@@
  • ABC Dataset: The ABC Dataset (Koch et al., 2019) is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications. It is currently one of the largest publicly available datasets for geometric deep learning research.
@@
Training details: Straightforwardly, the PU-GCN used in this article, where the network structure is shown in the architecture of (a) in Figure  2, is implemented using PyTorch. Due to TensorFlow and CUDA versions, this article is unable to re-create the original author’s Python environment in PU-GCN within the TensorFlow deep learning framework. And the experiments were conducted on a single NVIDIA GeForce RTX 3060 LHR GPU running on a Ubuntu 22.04 operating system. In terms of training hyperparameters, this article follows the settings of the original PU-GCN in terms of network structure. Notwithstanding, a batch size of $32$ with $50$ training epochs was used, with an initial learning rate of $5\times 10^{-4}$, a decay rate of $0.05$, and so on.
@@
Figure 2. The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure. (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
3.2. Evaluation Metrics
@@
During the testing process, this article adopted the common practices used in previous work on point cloud upsampling (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022). That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points.
@@
In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed. Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points. This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) (Hajdu et al., 2012), (ii) Hausdorff distance (HD) (Berger et al., 2013), and (iii) model parameter quantity.
3.3. Results
@@
PU1K dataset: Table  1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset. By upsampling at resolutions of 256-1024 and 2048-8192, it can be seen that the Patch based method outperforms the AS method proposed in this paper in all two metrics of distance measurement.
@@
Table 1. The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
@@
ABC10K dataset: we also conducted point cloud upsampling experiments using the ABC10K dataset. Perform upsampling at a resolution of 512-2048. As shown in Table  2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
Table 2. The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
@@
Although article proposed name Average Segment (AS) method has many advantages in theory, experimental results show that the Patch based upsampling method performs better on Chamfer distance and Hausdorff distance. One possible reason for this may be that when upsampling on a patch, each patch contains fewer point clouds, and the graph convolutional network can extract local features more finely. Ergo, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
3.4. Ablations
Test sets cross validation: This article will conduct cross validation on the PU-GCN trained on the PU1K and ABC10K datasets under the validation sets of both datasets. To understand the impact of the model used for training data on upsampling results. The results are shown in Table  3. When the Patch method and AS method are trained on the PU1K dataset and tested on the ABC10K dataset, the CD and HD metrics are significantly superior to those trained and tested on the ABC10K dataset. However, when trained on the ABC10K dataset and tested on the PU1K dataset, there was no significant difference between the CD and HD indicators.
Table 3. The results of swapping test sets based on Patch method and AS method on two datasets.
@@ 3.4. Ablations
Architecture Impact: This article breaks down the important modules of PU-GCN, aiming to explore which part has the greatest impact on point cloud upsampling. To this end, the original network backbone has been increased or decreased in three ways: removing DenseGCN (b), adding Refiner (c), and removing DenseGCN while adding Refiner (d), all of them are showed in Figure  2. The experiment was trained on the PU1K dataset and tested on the test sets of the PU1K dataset and ABC10K dataset. The results are shown in Table  4. The results showed that upsampling on the test set with three resolutions of 256-1024, 512-2048 and 2048-8192 in PU1K performed better than the original PU-GCN with DenseGCN, regardless of the presence or absence of Refiner, under the condition of removing DenseGCN. Under the upsampling scale of 2048-8192, the uniform segmentation method that removes DenseGCN is better than the Patch method.
Table 4. Perform ablation experiments on different architectures of PU-GCN using two data input methods.
@@
w/o Inception DenseGCN Patch PU1K $256\to 1024$ 35.061 54.340 99.152
@@
w/ Refiner w/o Inception DenseGCN Patch PU1K $256\to 1024$ 34.309 54.070 130.945
@@
3.5. Visualizations
The qualitative results of upsampling using different point cloud datasets are shown in Figure  3 and Figure  4. It can be seen that the input based on Patch performs better in distinguishing edges on the boundaries compared to the AS method. This is particularly evident on the CAD based ABC dataset, as shown in Figure  4. However, in Figure  3, it can be seen that the AS method performs better locally in the model. Relatively speaking, it is denser and the connections between points are tighter. In addition, even though DenseGCN performs better on CD and HD on the PU1K dataset, this results in the point cloud lacking local features in visualization.
Figure 3. Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
Figure 4. Visualization example on the ABC dataset.
4. Discussion